# Nucleon energy correlators as a probe of light-quark dipole operators at the EIC

### v1

@@ Abstract
Conclusion and discussion In summary, we have shown that nucleon energy correlators (NECs) provide a powerful framework to probe EW light-quark dipole operators in inclusive DIS with an unpolarized nucleon. By exploiting azimuthal correlations of the energy flux in the TFR, we introduce a chiral-odd NEC sensitive to the transverse spin of quarks, even within an unpolarized target. Coupling to the interference between the dipole and SM amplitudes, the transversity NEC produces characteristic $\sin\phi$ and $\cos\phi$ asymmetries in the energy pattern, providing clean access to the dipole couplings. Our combined analysis with both unpolarized and longitudinally polarized electron beams demonstrates that these observables can deliver precise and complementary constraints at the EIC.
@@
This work initiates an exploration of energy-correlator observables in the search for new physics (for an effort in other directions, see Ricci:2022htc ). Compared with proposals based on transversity PDFs or dihadron production Boughezal:2023ooo ; Wen:2024cfu , our NEC-based approach eliminates the need for nucleon polarization and relies entirely on inclusive calorimetric measurements, without particle identification or hadron reconstruction. These features simplify experimental implementation and reduce statistical uncertainties. They also enable the full utilization of high luminosity from the future EIC and LHeC as well as the use of existing HERA data. Furthermore, the resulting constraints might benefit from extensions to heavy-ion beams.
@@
While the present study employs inclusive NECs, improved angular resolution could be achieved through track-based measurements of charged-hadron energy flux Chen:2020vvp ; Li:2021zcf ; Jaarsma:2023ell ; Lee:2023npz , which would be valuable for precision studies on the dipole couplings. Furthermore, the chiral-odd NEC framework can be extended to polarized nucleons, with all such NECs sharing identical hard-scattering kernels, thus offering robust and complementary constraints. Beyond DIS, NECs could also be applied to processes such as Drell–Yan production at the LHC. Detailed investigations along these lines are left for future work.
@@ Acknowledgements.
Acknowledgments We thank Yu Jia, Yi Liao, Xiaohui Liu, Jianping Ma, Hongxi Xing for helpful discussions and comments. X.B. Tong is supported by the Research Council of Finland, the Centre of Excellence in Quark Matter and projects 338263 and 359902, and by the European Research Council (ERC, grant agreements No. ERC-2023-101123801 GlueSatLight and No. ERC-2018-ADG-835105 YoctoLHC). H. L. Wang is supported by the Grants No. NSFC-12247151. The content of this article does not reflect the official opinion of the European Union and responsibility for the information and views expressed therein lies entirely with the authors.
@@
Appendix A The complete results for energy-pattern cross section in the SM and SMEFT
@@
Here we summarize the complete results for the energy-pattern cross section in the SM and SMEFT. In this work, we consider the case, where the target nucleon is unpolarized and the electron beam is unpolarized or longitudinally polarized. We are interested in the following components:
@@
For illustration, Fig. 4 shows the numerical results for $|A_{UU}^{\sin\phi}|$ and $|A_{LU}^{\sin\phi}|$ as a function of the Bjorken variable $x_{B}$ for the up-quark dipole operators with different $Q$ at the EIC kinematics. We take $\sqrt{s}=105\ \mathrm{GeV}$ to optimize luminosity and apply an inelasticity cut of $0.1<y<0.9$ AbdulKhalek:2021gbh . The dimensionful dipole couplings are fixed to $c_{u\gamma}=v/(\sqrt{2}\,\mathrm{TeV}^{2})$ for both asymmetries, with $c_{uZ}=v/(\sqrt{2}\,\mathrm{TeV}^{2})$ for $|A_{UU}^{\sin\phi}|$ and $c_{uZ}=200\,v/(\sqrt{2}\,\mathrm{TeV}^{2})$ for $|A_{LU}^{\sin\phi}|$ to enhance visibility. We find $A_{UU}^{\sin\phi}>0$ for $c_{u\gamma}$ and $A_{UU}^{\sin\phi}<0$ for $c_{uZ}$, while $A_{LU}^{\sin\phi}<0$ for $c_{u\gamma}$ and $A_{LU}^{\sin\phi}>0$ for $c_{uZ}$. The magnitude of both asymmetries grows with $Q$. In particular, $A_{LU}^{\sin\phi}$ from $c_{u\gamma}$ is free from the $Q^{2}/m_{Z}^{2}$ suppression, almost reaching $\mathcal{O}(10^{-2})$ at $Q=30$ GeV. Results for $d$-quark operators follow a similar pattern, with slightly reduced magnitudes due to the smaller $h_{1}^{t,d}$. In our analysis, we use TMDs as the input for the NECs and apply the parameterization from Barone:2009hw . The resulting asymmetries are consistent in magnitude with those obtained using the alternative fit in Lu:2009ip .
Appendix B Additional constraints on the dipole operators at the EIC

### v2

@@ Abstract
Conclusion and discussion In summary, we have shown that nucleon energy correlators provide a powerful framework to probe EW light-quark dipole operators in inclusive DIS with an unpolarized nucleon. By exploiting azimuthal correlations of the energy flux in the TFR, we introduce a chiral-odd NEC sensitive to the transverse spin of quarks, even within an unpolarized target. Coupling to the interference between the dipole and SM amplitudes, the transversity NEC produces characteristic $\sin\phi$ and $\cos\phi$ asymmetries in the energy pattern, providing clean access to the dipole couplings. Our combined analysis with both unpolarized and longitudinally polarized electron beams demonstrates that these observables can deliver precise and complementary constraints at the EIC.
@@
This work initiates an exploration of energy-correlator observables in the search for new physics (for an effort in other directions, see Ricci and Riembau (2022)). Compared with proposals based on transversity PDFs or dihadron production Boughezal et al. (2023); Wen et al. (2024a), our NEC-based approach eliminates the need for nucleon polarization and relies entirely on inclusive calorimetric measurements, without particle identification or hadron reconstruction. These features simplify experimental implementation and reduce statistical uncertainties. They also enable the full utilization of high luminosity from the future EIC and LHeC as well as the use of existing HERA data. Furthermore, the resulting constraints may benefit from extensions to heavy-ion beams.
@@
While the present study employs inclusive NECs, improved angular resolution could be achieved through track-based measurements of charged-hadron energy flux Chen et al. (2020); Li et al. (2022); Jaarsma et al. (2023); Lee and Moult (2023), which would be valuable for precision studies on the dipole couplings. Furthermore, the chiral-odd NEC framework can be extended to polarized nucleons, with all such NECs sharing identical hard-scattering kernels, thus offering robust and complementary constraints. Beyond DIS, NECs could also be applied to processes such as Drell–Yan production at the LHC. Detailed investigations along these lines are left for future work.
@@ Acknowledgements.
Acknowledgments We thank Yu Jia, Yi Liao, Xiaohui Liu, Jianping Ma, Hongxi Xing for helpful discussions and comments. X.B. Tong is supported by the Research Council of Finland, the Centre of Excellence in Quark Matter and projects 338263 and 359902, and by the European Research Council (ERC, grant agreements No. ERC-2023-101123801 GlueSatLight and No. ERC-2018-ADG-835105 YoctoLHC). H. L. Wang is supported by the Grants No. NSFC-12035008. The content of this article does not reflect the official opinion of the European Union and responsibility for the information and views expressed therein lies entirely with the authors.Note added On the day this paper was submitted to arXiv, a preprint Michel:2025afc appeared in which the author proposed using a chiral-odd fracture function to probe Yukawa couplings at the LHC. In contrast to NECs, fracture functions typically involve measuring a specific type of hadron with momentum $\vec{P}_{h}$ in the TFR (see e.g., Anselmino:2011ss; Chen:2023wsi; Chen:2024brp; Chen:2021vby; Chai:2019ykk), which requires particle identification and tracking. While Michel:2025afc focused on the $\vec{P}_{h}$-integrated fracture functions, we note that the $\vec{P}_{h}$-dependent fracture functions (also referred to as extended fracture functions Grazzini:1997ih) are formally related to NECs through an inclusive energy sum rule recently established in Chen et al. (2024).
@@
Appendix A Quark transversity NEC
In Eq. (4) of our manuscript, we introduced the quark transversity NEC by definition. Here we provide additional details on its derivation and physical interpretation. We begin with the general quark NEC correlation matrix,
${\cal M}_{ij}(x,\vec{v})$ $=\int\frac{d\eta^{-}}{2\pi}e^{-ixP^{+}\eta^{-}}\langle P|\bar{\psi}_{j}(\eta^{-}){\cal E}(\vec{v})\psi_{i}(0)|P\rangle~.$ (14)
where the three-vector $\vec{v}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$ specifies the spatial direction of the measured energy flux. For brevity, we have suppressed the gauge links ${\cal L}_{n}$.
In analogy with the decomposition of collinear and TMD parton distributions Jaffe (1996); Boussarie and others (2023), hermiticity and parity constrain the NEC correlator to three independent Dirac structures at leading twist:
${\cal M}_{ij}(x,\vec{v})=\frac{1}{2}\bigg[(\gamma^{-})_{ij}{\cal M}^{[\gamma^{+}]}+(\gamma_{5}\gamma^{-})_{ij}{\cal M}^{[\gamma^{+}\gamma_{5}]}-i\frac{(\sigma_{\alpha_{\perp}}^{~-}\gamma_{5})_{ij}}{2}{\cal M}^{[i\sigma^{\alpha_{\perp}+}\gamma_{5}]}\bigg]+{\cal O}\Big(\frac{M}{P^{+}}\Big)~,$ (15)
where each component is obtained through the projection $\mathcal{M}^{[\Gamma]}=\frac{1}{2}\operatorname{Tr}[\Gamma\mathcal{M}]$. Because all leading-twist projections involve $\gamma^{+}$, they receive the dominant contributions in the high-energy limit ($P^{+}\rightarrow\infty$). Here, the chiral-even components ${\cal M}^{[\gamma^{+}]}$ and ${\cal M}^{[\gamma^{+}\gamma_{5}]}$, which describe the distributions of unpolarized or longitudinally polarized quarks, were discussed in detail in Ref. Chen et al. (2024) for both unpolarized and polarized targets.
We focus on the chiral-odd component ${\cal M}^{[i\sigma^{\alpha_{\perp}+}\gamma_{5}]}$ , where $\Gamma^{\alpha}_{\perp}=i\sigma^{\alpha_{\perp}+}\gamma_{5}$ with $\alpha_{\perp}=1,2$ is the projector for quark transverse spin:
${\cal M}^{[i\sigma^{\alpha_{\perp}+}\gamma_{5}]}(x,\vec{v})=\int\frac{d\eta^{-}}{2\pi}e^{-ixP^{+}\eta^{-}}\langle P|\bar{\psi}(\eta^{-}){\cal E}(\vec{v})i\sigma^{\alpha_{\perp}+}\gamma_{5}\psi(0)|P\rangle~.$ (16)
The interpretation of quark transverse spin becomes explicit by rewriting
$\Gamma^{\alpha}_{\perp}=\gamma^{+}({\cal Q}_{+}^{\alpha}-{\cal Q}_{-}^{\alpha})~,$ (17)
where ${\cal Q}_{\pm}^{\alpha}=\frac{1}{2}\big(1\mp\gamma^{5}\gamma^{\alpha}_{\perp}\big)$ projects onto quarks with transverse spin aligned or anti-aligned with the $\alpha_{\perp}$-direction Jaffe (1996); Jaffe and Ji (1992). Because transverse quark spin is invariant under longitudinal boosts, the corresponding NEC component ${\cal M}^{[i\sigma^{\alpha_{\perp}+}\gamma_{5}]}$ must transform covariantly under transverse rotations. For an unpolarized target, the only available transverse vector in the correlator is the transverse direction of the energy flow, ${\hat{n}}_{T}^{\alpha}\equiv\frac{{\bm{v}}_{\perp}^{\alpha}}{|{\bm{v}}_{\perp}|}=(\cos\phi,\sin\phi)$, so the dependence on the transverse index $\alpha_{\perp}$ must appear through the unique pseudovector structure
$\mathcal{M}^{\left[i\sigma^{\alpha_{\perp}+}\gamma_{5}\right]}(x,\vec{v})=\epsilon_{\perp}^{\alpha\beta}\hat{n}_{T,\beta}h^{t}_{1}(x,\theta^{2})~,$ (18)
where the transverse Levi-Civita tensor $\epsilon_{\perp}^{\alpha\beta}$ appears because the quark spin is a pseudovector. Eq. (18) implies a $\sin\phi$ or $\cos\phi$ azimuthal correlation between the quark’s transverse polarization and the azimuthal direction of the energy flow.
The non-perturbative function $h_{1}^{t}(x,\theta^{2})$, which we refer to as the quark transversity NEC, characterizes the strength of this azimuthal correlation as well as its dependence on $x$ and $\theta$. Using Eq. (17), it can be explicitly written as:
$h^{t}_{1}(x,\theta^{2})=\int\frac{d\eta^{-}}{4\pi}e^{-ixP^{+}\eta^{-}}\langle P|\bar{\psi}(\eta^{-}){\cal E}(\vec{v})\gamma^{+}\epsilon^{\perp}_{\alpha\beta}\hat{n}_{T\perp}^{\beta}({\cal Q}_{+}^{\alpha}-{\cal Q}_{-}^{\alpha})\psi(0)|P\rangle~.$ (19)
It describes the difference in probability of finding a quark whose transverse polarization is aligned with the pseudovector $\epsilon^{\perp}_{\alpha\beta}\hat{n}_{T\perp}^{\beta}$ versus the opposite orientation. Here, the pseudovector $\epsilon^{\perp}_{\alpha\beta}\hat{n}_{T\perp}^{\beta}$ is perpendicular to the transverse direction of the energy flow, $\hat{n}_{T}$. Thus, the quark transversity NEC $h_{1}^{t}(x,\theta^{2})$ characterizes how the transverse polarization of a quark inside an unpolarized nucleon correlates with the direction of the energy flow detected in the target-fragmentation region.
Appendix B Factorization of the DIS energy pattern and the dipole-induced azimuthal modulations in the TFR
In our manuscript, we show that the $\sin\phi$ and $\cos\phi$ modulations of the DIS energy pattern uniquely arise from the coupling of the quark transversity NEC to the dipole-SM interference, as presented in Eqs. (6) and (10) of the main body. Here we provide additional discussion of the derivation from the factorization point of view. First, this derivation follows the general factorization theorem for the DIS energy pattern in the target fragmentation region (TFR) within the SM Cao et al. (2023); Chen et al. (2024), with the additional inclusion of light-quark dipole interactions in the hard partonic scattering part $\mathcal{H}_{ij}$ and the quark transversity NEC in the nonperturbative parton correlation matrix $\mathcal{M}_{ij}$, as illustrated in Figure 4.
Figure 4: The factorized structure of the DIS energy pattern in the TFR with the inclusion of the light-quark dipole interactions. The blue blob denotes the dipole interactions.
At the leading order of $\alpha_{s}$, the factorized structure takes the form:
$\frac{d\Sigma(\theta,\phi)}{dx_{B}dQ^{2}}\propto\int_{x_{B}}^{1}\frac{dx}{x}{L}_{\mu\nu}(l,l^{\prime}){\cal H}_{ji}^{\mu\nu}\Big(\frac{x_{B}}{x}\Big)\delta(1-\frac{x_{B}}{x}){\cal M}_{ij}(x,\theta,\phi)~,$ (20)
where $L_{\mu\nu}$ denotes the leptonic tensor. This formula captures the leading-power contributions in the TFR, valid in the kinematic regime $\theta P^{+}\ll Q$ under the Bjorken limit $Q\gg\Lambda_{\mathrm{QCD}}$. In this region, the observed energy flow originates from the fragmentation of the target remnants after the virtual photon or $Z$-boson knocks out a quark from the proton. The factorization arises because the long-distance target–fragmentation dynamics is well separated from the short-distance hard scattering, but remains strongly correlated with the struck quark. This long-distance correlation is encoded in the NEC correlator $\mathcal{M}_{ij}$, whereas the hard scattering between the quark and the virtual probe, $\mathcal{H}_{ij}$ is perturbatively calculable. Moreover, since we observe only the energy flow in the TFR and not in the current-fragmentation region, the transverse momentum of the struck quark is not resolved. As a result, the quark $k_{\perp}$-dependence in $\mathcal{M}_{ij}$ integrates out, yielding a collinear factorization structure. As demonstrated in Ref. Chen et al. (2024), the TFR factorization formula for the DIS energy pattern can be rigorously derived from the TFR factorization theorem for single-hadron inclusive DIS, originally proven by Collins in a seminal work Collins (1998).
In the SM, the hard scattering part ${\cal H}^{\mu\nu}_{ij}$ preserves chirality for light quarks. Therefore, the chiral-odd component $\mathcal{M}^{\left[i\sigma^{\alpha}\perp^{+}\gamma_{5}\right]}$ in Eq. (15) does not contribute, and no azimuthal modulation arises for an unpolarized target. However, once the interference between the dipole operator and the SM amplitude is included, the hard part ${\cal H}^{\mu\nu}_{ij}$ contains an even number of $\gamma$-matrices, which allows a nonzero contraction with the chiral-odd structure:
$\frac{d\Sigma(\theta,\phi)}{dx_{B}dQ^{2}}\propto{L}_{\mu\nu}(l,l^{\prime})\text{Tr}\Big[{\cal H}^{\mu\nu}i\sigma^{-\alpha_{\perp}}\gamma_{5}\Big]\epsilon_{\perp}^{\alpha\beta}\hat{n}_{T,\beta}h^{t}_{1}(x_{B},\theta^{2})~.$ (21)
In our setup, the azimuthal angle $\phi$ of the energy flux is measured relative to the lepton scattering plane in the Breit frame. The transverse axes are chosen such that the leptons carry the transverse momentum $\bm{l}^{\mu}_{\perp}=-\bm{l}^{\prime\mu}_{\perp}=(Q\sqrt{1-y}/{y},0)$. Because the partonic hard kernel ${\cal H}^{\mu\nu}$ does not depend on any transverse momentum, by rotational covariance, the contraction of the leptonic tensor with ${\cal H}^{\mu\nu}$ must reduce to:
${L}_{\mu\nu}(l,l^{\prime})\text{Tr}\Big[{\cal H}^{\mu\nu}i\sigma^{-\alpha_{\perp}}\gamma_{5}\Big]={\cal C}_{1}l^{\alpha}_{\perp}+{\cal C}_{2}\epsilon_{\perp}^{\beta\alpha}l_{\perp,\beta}~.$ (22)
where $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are scalar functions of the invariants, linearly proportional to the real and imaginary parts of the quark dipole couplings, respectively. Contracting this structure with the NEC correlator yields the characteristic $\sin\phi$ and $\cos\phi$ modulations:
$\frac{d\Sigma(\theta,\phi)}{dx_{B}dQ^{2}}\propto\Big[{\cal C}_{1}\sin\phi+{\cal C}_{2}\cos\phi\Big]h^{t}_{1}(x_{B},\theta^{2})~.$ (23)
This demonstrates how the quark transversity NEC $h_{1}^{t}\left(x_{B},\theta^{2}\right)$ provides the probe for the dipole-SM interference and generates the unique azimuthal modulations of the DIS energy pattern in the TFR.
Appendix C Evolution of the NEC moments in phenomenological analysis
In this section, we describe how scale evolution is treated for NECs used in our phenomenological analysis. The relevant momentum scale is identified as the virtuality $Q$ in the DIS process. As discussed in the main text, the observables considered in this work are expressed in terms of $\theta$-weighted moments of NECs. These moments can be related to $\bm{k}_{\perp}$-moments of TMDs, which allows us to connect NECs to existing phenomenological fits of TMDs.
We adopt the TMD fits as the initial conditions for the NEC moments at the scale $Q_{0}=2\mathrm{GeV}$ and evolve them to other values of $Q$. This evolution is performed under the approximation that the $\theta$-moments of the NECs obey the same evolution equations as the corresponding NECs themselves.
We now justify this approximation by examining the evolution properties of the NECs. The unpolarized quark NEC $f_{1}\left(x,\theta^{2}\right)$ and the quark transversity NEC $h_{1}^{t}\left(x,\theta^{2}\right)$ obey the same renormalization group evolution equations as the unpolarized and transversity collinear quark PDFs, respectively. This follows from the fact that NECs can be formally expressed as collinear PDFs with insertions of energy-flow operators. Since the energy-flow operator acts on the inclusive final state, its insertion does not modify the ultraviolet structure of the operator and therefore does not affect the corresponding evolution equations Chen et al. (2024).
The observables considered in our analysis are related to $\theta$-moments of NECs. For quark transversity, the relevant moment is defined as
$H_{1}^{t}(x)\equiv E_{N}\int d\theta^{2}\,|\sin\theta|\,h_{1}^{t}(x,\theta^{2})=\int\frac{d^{2}\bm{k}_{\perp}}{2\pi}\,\frac{\bm{k}_{\perp}^{2}}{M}\,h_{1,\mathrm{TMD}}^{\perp}(x,\bm{k}_{\perp}^{2}).$ (24)
which can be equivalently written as a $\bm{k}_{\perp}$-moment of the Boer–Mulders TMD $h_{1,\mathrm{TMD}}^{\perp}$. In principle, the evolution of this NEC moment differs from that of the NEC itself due to the appearance of additional inhomogeneous terms. From the known properties of the $\bm{k}_{\perp}$-moment of the BM TMD, the moment $H_{1}^{t}(x,Q)$ is related to the twist-3 quark–gluon correlator $T_{q,F}^{(\sigma)}(x,x,Q)$,
$H_{1}^{t}(x,Q)\propto T_{q,F}^{(\sigma)}(x,x,Q),$ (25)
whose QCD evolution has been derived in Kang and Qiu (2012).
However, if one retains only the homogeneous term in the evolution equation, the moment $H_{1}^{t}(x,Q)$ evolves according to the same DGLAP equation as the transversity PDF Artru and Mekhfi (1990); Vogelsang (1998) and, correspondingly, the transversity NEC:
$\frac{\partial}{\partial\ln Q^{2}}H_{1}^{t}(x,Q)=\frac{\alpha_{s}}{2\pi}\int_{x}^{1}\frac{d\hat{x}}{\hat{x}}P_{q\to q}^{h_{1}}(\hat{x})\,H_{1}^{t}\!\left(\frac{x}{\hat{x}},Q\right),$ (26)
with the splitting kernel
[…]
In our phenomenological study, this homogeneous evolution is adopted as a reasonable approximation for estimating scale dependence. In other words, we assume the $\theta$-moments of the NECs follow the same evolution equations as the corresponding NECs themselves. Under this assumption, the evolution of the unpolarized NEC moment is automatically included through the collinear evolution encoded in the TMD fits.
@@ Appendix C Evolution of the NEC moments in phenomenological analysis
Appendix D The complete results for energy-pattern cross section in the SM and SMEFT
@@
Here we summarize the complete results for the energy-pattern cross section in the SM and SMEFT. In this work, we consider the case where the target nucleon is unpolarized and the electron beam is unpolarized or longitudinally polarized. We are interested in the following components:
@@
For illustration, Fig. 5 shows the numerical results for $|A_{UU}^{\sin\phi}|$ and $|A_{LU}^{\sin\phi}|$ as a function of the Bjorken variable $x_{B}$ for the up-quark dipole operators with different $Q$ at the EIC kinematics. We take $\sqrt{s}=105\ \mathrm{GeV}$ to optimize luminosity and apply an inelasticity cut of $0.1<y<0.9$ Abdul Khalek and others (2022). The dimensionful dipole couplings are fixed to $c_{u\gamma}=v/(\sqrt{2}\,\mathrm{TeV}^{2})$ for both asymmetries, with $c_{uZ}=v/(\sqrt{2}\,\mathrm{TeV}^{2})$ for $|A_{UU}^{\sin\phi}|$ and $c_{uZ}=200\,v/(\sqrt{2}\,\mathrm{TeV}^{2})$ for $|A_{LU}^{\sin\phi}|$ to enhance visibility. We find $A_{UU}^{\sin\phi}>0$ for $c_{u\gamma}$ and $A_{UU}^{\sin\phi}<0$ for $c_{uZ}$, while $A_{LU}^{\sin\phi}<0$ for $c_{u\gamma}$ and $A_{LU}^{\sin\phi}>0$ for $c_{uZ}$. The magnitude of both asymmetries grows with $Q$. In particular, $A_{LU}^{\sin\phi}$ from $c_{u\gamma}$ is free from the $Q^{2}/m_{Z}^{2}$ suppression, almost reaching $\mathcal{O}(10^{-2})$ at $Q=30$ GeV. Results for $d$-quark operators follow a similar pattern, with slightly reduced magnitudes due to the smaller $h_{1}^{t,d}$. In our analysis, we use TMDs as the input for the NECs and apply the parameterization from Barone et al. (2010). The resulting asymmetries are consistent in magnitude with those obtained using the alternative fit in Lu and Schmidt (2010).
Appendix E Additional constraints on the dipole operators at the EIC